prove that 1.2+2.2 2 +3.2 3 +............+100.2 100 =99.2 101 +2.

  1. prove that 1.2+2.22+3.23+............+100.2100=99.2101+2.

Grade:11

1 Answers

moumi roy
7 years ago
S=1.2+2.22+3.23+............+100.2100
2S=    +1.22+2.23+..........+99.2100+100.2101
SUBSTRACTING,S=100.2101-[2+22....+2100]=100.2101-2(2100-1 )=99.2101+2.